# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

### v1

@@ 4.3.3. Bound for $III$: Smooth minimum approximation
We study the term $III$ to evaluate the approximation error by the smooth minimum function $F_{\beta}$. A straightforward calculation yields approximation to minimum function as
@@
$\left|E\left[g\left(\min_{w\in\mathcal{W}}H_{\ell_{\rho}}(w,\mathbf{X})\right)\right]-E\left[g\left[n^{-1}F_{\beta}\left((\bar{H}_{\ell_{\rho}}(w,\mathbf{X}))_{w\in\mathcal{W}}\right)\right]\right]\right|\leq\frac{\|g^{{}^{\prime}}\|_{\infty}\log|W|}{n\beta}$
Evaluation by the soft minimum function is one of the useful tools used in high-dimensional analysis, as demonstrated by Chernozhukov et al. (2013). By exploiting the smoothness of the soft minimum function, we can apply various approximation methods, such as Stein’s method and the generalized Lindeberg principle.
@@
In this proof, we employ several techniques used in the high-dimensional Gaussian approximation studies (Chernozhukov et al., 2013, 2014; Chernozhuokov et al., 2022; Lopes, 2022; Chang et al., 2024). However, not all techniques from the approximation can be applied directly. There are two main reasons for this. First, there is a difference in the objects of differentiation when considering neighborhood approximations. In the approximation studies, one needs to differentiate the smoothly approximated maximum of a high-dimensional data vector with respect to the high-dimensional data itself. In contrast, in our setting, the loss function maximized with respect to the parameters is differentiated by the values of the high-dimensional data. Since the values being maximized and differentiated differ in this way, we cannot apply the approximation studies in our setup. Second, the use of Nazarov’s inequality plays a crucial role. While some of the approximation study Chang et al. (2024) needs Nazarov’s inequality, Nazarov’s inequality also does not hold for the design matrix $\mathbf{X}$ or $\mathbf{W}$, since we consider the explicit form of the minimum of the empirical risk function. Due to these differences, the results from the studies on high-dimensional Gaussian approximations cannot be directly applied.
5. Application to Distribution of Robust Estimators
@@
In this section, we demonstrate the universality of the distribution of estimators with block dependence under specific loss functions designed for robustness. Specifically, we consider estimators under the absolute loss and the Huber loss, as presented in Example 1. We define a robust cost function as
$H^{R}(w,\mathbf{X},\xi):=\frac{1}{n}\sum_{i=1}^{n}\ell_{0}((X_{i}^{\top}w)_{i}-\xi_{i})+\frac{\lambda}{2}\|w+\theta_{0}\|^{2}.$
@@
where $\lambda>0$ is a coefficient for the regularizer. Let $\widehat{w}^{R}_{\mathbf{X}}$ denote the minimizer of the cost function with covariates $\mathbf{X}$. Recall that $\theta_{0}\in\mathbb{R}^{p}$ be the true parameter of the regression model (1), and the estimator of the regression model as the minimizer (2) with the setup is given as $\widehat{\theta}_{\mathbf{X}}^{R}=\widehat{w}_{\mathbf{X}}^{R}+\theta_{0}$. By substituting a Gaussian design matrix $\mathbf{G}$ instead of $\mathbf{X}$, we can consider the Gaussian analogue $H^{R}(w,\mathbf{G},\xi)$ as Section 3.1.
@@
Assumption 4 (Assumptions on Robust estimators).
The followings hold:
  1. i.
@@
  3. iii.
    A rescaled covariate $X^{0}_{i}=\sqrt{n}X_{i}$ satisfies $M_{k^{*}:\mathbf{X}}:=\max_{j\in[p]}E[|X^{0}_{i,j}|^{2\lceil k^{*}/2\rceil}]<\infty$ where $k^{*}>{3}/({2\varepsilon})$. Also, $\lambda_{\max}(\Sigma)<\infty$ holds.
Condition i generalizes the condition previously mentioned for the robustness of the loss. Both the absolute loss and the Huber loss satisfy this condition. By combining Conditions ii and iii, we can explain how the moment condition of the scaled covariates $X_{i}^{0}$ is affected by the order of $d$ for the block dependence.
@@
Suppose Assumptions 1 and 4 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Then, for any $k\geq 2$, we have $K:=K(p)>0$ satisfying
@@
Suppose Assumptions 1 and 4 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Fix a measurable subset $\mathcal{S}_{n}\subset\mathbb{R}^{p}$. Suppose that there exist $z\in\mathbb{R},\rho_{0}>0$, and $\varepsilon_{n}\in[0,1/4)$ such that the following hold:
  1. (ii)
@@
Then, $\widehat{w}_{\mathbf{X}}$ also does not satisfy the specific structure $\mathcal{S}_{n}$ with high probability, i.e. there exists some $K=K(\lambda,\tau,M_{k^{*}:\mathbf{X}},k^{*},\varepsilon,L_{0})>0$ such that
$\mathbb{P}(\widehat{w}^{R}_{\mathbf{X}}\in\mathcal{S}_{n})\leq 4\varepsilon_{n}+K(1+\|\theta_{0}\|^{k^{*}}_{\infty}+\rho^{-3}_{0})n^{-{1}/{45}}$
@@
With the result the universality of the distribution, Theorem 11 will derive the convergence result of estimation error with non-Gaussian covariates:
Theorem 12.
Suppose that Assumption 1 and 4 hold with $\sup_{p}\max_{j\in[p]}E[|X^{0}_{i,j}|^{2\lceil k^{*}/2\rceil}]<\infty$, and define $\tau_{0}=(n/p)$. Moreover. we assume the followings:
  1. (i)
    Either (a) $\ell_{0}$ does not have derivative at certain point, or (b) $\ell_{0}$ includes an interval where $\psi_{0}$ has a strictly increasing derivative.
  2. (ii)
  3. (iii)
    $\theta_{0}$ contains i.i.d. components whose law $\Pi_{0}$ guarantees the existence of moment of any order.
Then with $Z\sim N(0,1)$, the system of equations
$(\gamma^{R}_{*})^{2}/\tau_{0}=$ $E\left[\gamma^{R}_{*}Z+\xi_{1}-\mathsf{prox}_{\ell_{0}}\left(\gamma^{R}_{*}Z+\xi_{i};\beta^{R}_{*}\right)\right]+\lambda^{2}(\beta^{R}_{*})^{2}\cdot E\Pi^{2}_{0},$
$1-\tau_{0}^{-1}+\lambda\beta^{R}_{*}=$ $E\left[\mathbf{\mathsf{prox}^{\prime}}_{\ell_{0}}\left(\gamma^{R}_{*}Z+\xi_{i};\beta^{R}_{*}\right)\right]$
have a unique solution $(\beta^{R}_{*},\gamma^{R}_{*})\in(0,\infty)$ such that we have
[…]
@@
This result demonstrates that the error of the estimator $\widehat{\theta}_{\mathbf{X}}$ form non-Gaussian covariates converges to the known value $\tau_{0}(\gamma^{R}_{*})^{2}$, which is given by the solution of the system. The specification of the system has been well studied in the context of Gaussian covariates, e.g. Thrampoulidis et al. (2015, 2018). Our result shows that this universality of the estimation error also holds when the independence assumption of covariates $X_{i}$ is relaxed.
Remark 3 (Possibility of extension to lasso/ridge regression).
@@
We discuss extending this result to the Lasso and Ridge regressions under the squared loss, as examined in Han and Shen (2023). A key point here is whether the upper bound for the estimator holds, as in Lemma 10 in the setup in this section. In Han and Shen (2023), the upper bound is derived under the assumption that each element of $X_{i}$ is independent, so the bound does not apply to our case directly. Instead, we can utilize the upper bound under dependent covariates, as in Lemma 7.3 of Lahiry and Sur (2024), which allows us to show the universality of the Lasso and Ridge regressions.

### v2

@@ 4.3.3. Bound for $III$: smooth minimum approximation
We study the term $III$ to evaluate the approximation error by the smooth minimum function $F_{\beta}$. A straightforward calculation yields an approximation to the minimum function as
@@
$\left|E\left[g\left(\min_{w\in\mathcal{W}}H_{\ell_{\rho}}(w,\mathbf{X})\right)\right]-E\left[g\left[n^{-1}F_{\beta}\left((\bar{H}_{\ell_{\rho}}(w,\mathbf{X}))_{w\in\mathcal{W}}\right)\right]\right]\right|\leq\frac{\|g^{{}^{\prime}}\|_{\infty}\log|W|}{n\beta}.$
Evaluation by the soft minimum function is a useful tool used in high-dimensional analysis, as demonstrated by Chernozhukov et al. (2013). By exploiting the smoothness of the soft minimum function, we can apply various approximation methods, such as Stein’s method and the generalized Lindeberg principle.
@@
In this proof, we employ several techniques used in the high-dimensional Gaussian approximation studies (Chernozhukov et al., 2013, 2014, 2022; Lopes, 2022; Chang et al., 2024). However, not all techniques from the approximation can be applied directly. There are two main reasons for this. First, there is a difference in the objects of differentiation when considering neighborhood approximations. In the approximation studies, one needs to differentiate the smoothly approximated maximum of a high-dimensional data vector with respect to the high-dimensional data itself. In contrast, in our setting, the loss function minimized with respect to the parameters is differentiated by the values of the high-dimensional data. Since the values being maximized and differentiated differ in this way, we cannot apply the approximation studies in our setup. Second, the use of Nazarov’s inequality plays a crucial role. While some approximation studies, such as Chang et al. (2024), need Nazarov’s inequality, Nazarov’s inequality also does not hold for the design matrix $\mathbf{X}$ or $\mathbf{W}$, since we consider the explicit form of the minimum of the empirical risk function. Due to these differences, the results from the studies on high-dimensional Gaussian approximations cannot be directly applied.
5. Application
In this section, we demonstrate the universality of the distribution of estimators with block dependence. First, we examine the asymptotic behavior of M-estimators under Gaussian covariates with a general covariance matrix. We prove a convergence result for M-estimators by generalizing the main theorem of Thrampoulidis et al. (2018). Then, we apply this extension to estimators with block dependence under specific loss functions.
5.1. Precise error of regularized M-estimators with general covariance
Consider the following regularization problem
$\widehat{\theta}\in\operatornamewithlimits{argmin}_{\theta\in\mathbb{R}^{p}}\left\{\mathcal{L}(\mathbf{y}-\mathbf{G}\theta)+\lambda f(\theta)\right\}\quad\text{where}\quad Y_{i}=G_{i}^{\top}\theta_{0}+\xi_{i}$ (15)
where we define $\mathcal{L}(\mathbf{y}-\mathbf{G}\theta)=\sum_{i=1}^{n}\ell_{0}(Y_{i}-G_{i}^{\top}\theta)$ where $G_{i}\sim\mathcal{N}(0,\Sigma)$ and $\xi_{i}$ is an independent noise variable. We consider the asymptotic behavior of M-estimators derived from equation (15). To establish the result, we assume the following:
Assumption 4.
There exists $a>0$ such that for any $n$, we have $a\leq\lambda_{min}(\Sigma)$.
Assumption 5.
(Summary functionals $L(\cdot)$ and $F(\cdot)$) The following conditions hold:
  1. (a)
    For any $c\in\mathbb{R}$ and $\tau>0$, we have continuous functions $L:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that
    $\frac{1}{n}\{e_{\mathcal{L}}(c\mathbf{g}+\xi;\tau)-\mathcal{L}(\mathbf{\xi})\}\xrightarrow[]{p}L(c,\tau).$
  2. (b)
    For any $c\in\mathbb{R}$ and $\tau>0$, we have continuous functions $F:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that we have
    $\frac{1}{p}\{e_{g}(c\mathbf{h}+\Sigma^{1/2}\theta_{0};\tau)-g(\Sigma^{1/2}\theta_{0})\}\xrightarrow[]{p}F(c,\tau),$
    where we define $g(\mathbf{x})$ as $f(\Sigma^{-1/2}\mathbf{x})$ and $\mathbf{g}$ and $\mathbf{h}$ follow normal distributions, $\mathcal{N}(0,I_{n})$ and $\mathcal{N}(0,I_{p})$, respectively. Note that $g(\Sigma^{1/2}\theta_{0})$ is equal to $f(\theta_{0})$.
  3. (c)
    At least one of the following holds. There exists constant $C>0$ such that $\|\xi\|_{2}/\sqrt{p}\leq C$ with probability approaching 1 (w.p.a.1), or, $\sup_{\mathbf{v}\in\mathbb{R}^{p}}\sup_{\mathbf{s}\in\partial\mathcal{L}(\mathbf{v})}\|\mathbf{s}\|_{2}<\infty$ for any $p\in\mathbb{N}$.
Assumption 6.
(Properties of $L(\cdot)$ and $F(\cdot)$, Assumption 2 of Thrampoulidis et al. (2018)) We say that Assumption 6 holds if all the following are true.
  1. (a)
    $\lim_{\tau\rightarrow 0^{+}}F(\tau,\tau)=0$ and $\lim_{c\rightarrow+\infty}\left\{\frac{c^{2}}{2\tau}-F(c,\tau)\right\}=+\infty$ for all $\tau>0$.
  2. (b)
    $\lim_{\tau\rightarrow 0^{+}}L(\alpha,\tau)<+\infty,$ $\lim_{\tau\rightarrow 0^{+}}L(0,\tau)=0$ and $\infty<L_{2,+}(0,0):=\lim_{\tau\rightarrow 0^{+}}L_{2,+}(0,\tau)\leq 0$.
  3. (c)
    $\frac{1}{n}\mathcal{L}(\xi)\xrightarrow[]{p}L_{0}\in[0,\infty]$. Also, $L_{0}=-\lim_{\tau\rightarrow+\infty}L(c,\tau)\geq-L(c,\tau^{\prime})$ for all $c\in\mathbb{R},\tau^{\prime}>0$.
  4. (d)
    If $L_{0}=+\infty$, then $\lim_{\tau\rightarrow+\infty}L(c,\tau)/\tau=0$ for all $c\in\mathbb{R}$.
These assumptions are extensions of those of Thrampoulidis et al. (2018) under a general covariance $\Sigma$. Specifically, Assumptions 5 and 6 are identical to Assumption 1 and Assumption 2 in Thrampoulidis et al. (2018) respectively, except for the part on $F$. Because of the existence of $\Sigma$, we need to rotate the true parameter $\theta_{0}$ and a regularized function $f$. The rotation alters the structure of $F$, resulting in a different asymptotic behavior of the error $\frac{1}{p}\|\widehat{\theta}-\theta_{0}\|^{2}_{\Sigma}$ relative to that in Thrampoulidis et al. (2018).
Under Assumptions 3 to 6, we obtain the generalization of Theorem 3.1 in Thrampoulidis et al. (2018).
Theorem 10.
Let Assumptions 3 to 6 hold and define $\tau_{0}=\lim_{n\rightarrow\infty}(n/p)$. If the following convex-concave minimax scalar optimization
$\inf_{\begin{subarray}{c}\alpha\geq 0\\ \tau_{g}>0\end{subarray}}\sup_{\begin{subarray}{c}\beta\geq 0\\ \tau_{h}>0\end{subarray}}\mathcal{\mathcal{D}}(\alpha,\tau_{g},\beta,\tau_{h})\\ :=\frac{\beta\tau_{g}}{2}+\tau_{0}L\left(\alpha,\frac{\tau_{g}}{\beta}\right)-\frac{\alpha\tau_{h}}{2}-\frac{\beta^{2}\alpha}{2\tau_{h}}+\lambda F\left(\frac{\alpha\beta}{\tau_{h}},\frac{\alpha\lambda}{\tau_{h}}\right)$ (16)
has a unique minimizer $\alpha_{*}$, then we have
$\frac{1}{p}\|\widehat{\theta}-\theta_{0}\|^{2}_{\Sigma}\xrightarrow[]{p}\alpha^{2}_{*}.$ (17)
This theorem is a generalization of Theorem 3.1 in Thrampoulidis et al. (2018), extending the covariance matrix from the identity matrix to a general matrix $\Sigma$. Especially, the result of Theorem 10 coincides with that of Theorem 3.1 in Thrampoulidis et al. (2018) when $\Sigma=I$ holds. Specifically, a difference is that the norm $\|\cdot\|_{\Sigma}$ used in (17) is weighted by $\Sigma$, and the function $F(\cdot)$ in (16) corresponds to the general $\Sigma$, as introduced in Assumption 5.
5.2. Application to estimators
@@
In this section, we demonstrate the universality of the distribution of estimators with block dependence under specific loss functions.
First, we consider estimators under the absolute loss and the Huber loss, as presented in Example 1. We define a robust cost function as
$H^{R}(w,\mathbf{X},\xi):=\frac{1}{n}\sum_{i=1}^{n}\ell_{0}((X_{i}^{\top}w)-\xi_{i})+\frac{\lambda}{2}\|w+\theta_{0}\|^{2}.$
@@
where $\lambda>0$ is a coefficient for the regularizer. Let $\widehat{w}^{R}_{\mathbf{X}}$ denote the minimizer of the cost function with covariates $\mathbf{X}$. Recall that $\theta_{0}\in\mathbb{R}^{p}$ is the true parameter of the regression model (1), and the estimator of the regression model as the minimizer (2) with the setup is given as $\widehat{\theta}_{\mathbf{X}}^{R}=\widehat{w}_{\mathbf{X}}^{R}+\theta_{0}$. By substituting a Gaussian design matrix $\mathbf{G}$ instead of $\mathbf{X}$, we can consider the Gaussian analogue $H^{R}(w,\mathbf{G},\xi)$ as in Section 3.1.
@@
Assumption 7 (Assumptions for estimators).
The following hold:
  1. (a)
@@
  3. (c)
    Uniformly in $p$, a rescaled covariate $X^{0}_{i}=\sqrt{n}X_{i}$ satisfies $M_{k^{*}:\mathbf{X}}:=\max_{j\in[p]}E[|X^{0}_{i,j}|^{2\lceil k^{*}/2\rceil}]$ $\infty$ where $k^{*}>{3}/({2\varepsilon})$. Also, $\lambda_{\max}(\Sigma)<\infty$ holds.
  4. (d)
  5. (e)
    $\theta_{0}$ contains i.i.d. components whose law $\Pi_{0}$ guarantees the existence of moments of any order.
In Assumption 7, the condition (a) generalizes the condition previously mentioned for the robustness of the loss. Both the absolute loss and the Huber loss satisfy this condition. By combining the conditions (b) and (c), we can explain how the moment condition of the scaled covariates $X_{i}^{0}$ is affected by the order of $d$ for the block dependence.
@@
Suppose Assumptions 1 and 7 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Then, for any $k\geq 2$, we have $K:=K(k)>0$ satisfying
@@
Suppose Assumptions 1 and 7 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Fix a measurable subset $\mathcal{S}_{n}\subset\mathbb{R}^{p}$. Suppose that there exist $z\in\mathbb{R},\rho_{0}>0$, and $\varepsilon_{n}\in[0,1/4)$ such that $\widehat{w}_{\mathbf{G}}$ does not satisfy the specific structure $\mathcal{S}_{n}$ in the sense that
@@
Then, $\widehat{w}_{\mathbf{X}}$ also does not satisfy the specific structure $\mathcal{S}_{n}$ with high probability, i.e. there exists some $K=K(\lambda,\tau,M_{k^{*}:\mathbf{X}},k^{*},\varepsilon,L_{0})>0$ such that it holds that
$\mathbb{P}(\widehat{w}^{R}_{\mathbf{X}}\in\mathcal{S}_{n})\leq 4\varepsilon_{n}+K(1+\|\theta_{0}\|^{k^{*}}_{\infty}+\rho^{-3}_{0})n^{-{1}/{45}}.$
@@
With Theorem 3 for the universality, Theorem 10 and Theorem 12 derive the convergence result of estimation error with non-Gaussian covariates as follows:
Theorem 13 (Robust estimators).
Suppose that Assumptions 1, 4, 5-(b) 6-(a), and 7 hold. Define $\tau_{0}=\lim_{n\rightarrow\infty}(n/p)$. Moreover, we assume either (a) $\ell_{0}$ does not have a derivative at a certain point, or (b) $\ell_{0}$ includes an interval where $\psi_{0}$ has a strictly increasing derivative. Then, if convex-concave minimax scalar optimization (16) has a unique minimizer $\alpha^{*}$, we have
$p^{-1}{\|\widehat{\theta}^{R}_{\mathbf{X}}-\theta_{0}\|_{\Sigma}^{2}}\overset{p}{\rightarrow}\alpha_{*}^{2}.$
@@
This result demonstrates that the error of the estimator $\widehat{\theta}^{R}_{\mathbf{X}}$ from non-Gaussian covariates converges to the known value $\alpha^{2}_{*}$, which is given as the solution of the convex-concave minimax scalar optimization defined in (16). Our result shows that this universality of the estimation error also holds when the independence assumption of covariates $X_{i}$ is relaxed.
Second, we apply our main result to the Lasso and Ridge regressions under the squared loss, as examined in Han and Shen (2023). Let $\widehat{w}^{Ri}$ and $\widehat{w}^{L}$ denote Ridge and Lasso estimators, respectively. Moreover, define $f^{Ri}$ and $f^{L}$ denote $\|\cdot\|_{2}$ and $\|\cdot\|_{1}$, respectively.
Theorem 14 (Ridge/ Lasso estimator).
Suppose that Assumptions 1, 3, 4, 5-(b), 6-(a), and 7-(d)(e) hold and define $\tau_{0}=\lim_{n\rightarrow\infty}(n/p)$. Moreover, we assume the second moment of $\xi$ is finite, $d$ increases so that $L_{n}^{7}(\log n)^{2}\sigma_{n}=o(1)$ holds as $n\to\infty$, and there exist $L_{n}\geq 1$ and $\varepsilon_{n}\in[0,1/4)$ such that both $\|\widehat{w}^{A}_{\mathbf{G}}\|_{\infty}$ and $\|\widehat{w}^{A}_{\mathbf{X}}\|_{\infty}$ increase moderately in the sense that for each $A\in\{Ri,L\}$, we have
$\max\left\{\mathbb{P}(\|\widehat{w}^{A}_{\mathbf{G}}\|_{\infty}>L_{n}),\mathbb{P}(\|\widehat{w}^{A}_{\mathbf{X}}\|_{\infty}>L_{n})\right\}\leq\varepsilon_{n}.$
Let $\alpha^{Ri}_{*}$ and $\alpha^{L}_{*}$ denote the unique minimizers of convex-concave minimax scalar optimization (16) with $F^{Ri}$ and $F^{L}$, respectively. Then, for each $A\in\{Ri,L\}$, we have
$p^{-1}{\|\widehat{\theta}^{A}_{\mathbf{X}}-\theta_{0}\|_{\Sigma}^{2}}\overset{p}{\rightarrow}(\alpha_{*}^{A})^{2}.$
@@
A key point here is whether the upper bound for the estimator holds, as in Lemma 11 in the setup in this section. In Han and Shen (2023), the upper bound is derived under the assumption that each element of $X_{i}$ is independent, so the bound does not apply directly to our case. Instead, we can utilize the upper bound under dependent covariates, as in Lemma 7.3 of Lahiry and Sur (2024), which allows us to show the universality of the Lasso and Ridge regressions.
6. Numerical experiments
We numerically investigate the universality of the distribution of the estimation error under dependent data. In particular, the cell size $d$ has the same order as the parameter dimension $p$.
We generate synthetic data from the linear regression model (1). To introduce dependence between the features, we partition the $p$ elements of inputs into cells of size $d$. Each block is generated independently, but features within each cell have a correlation $\rho=0.5$. Formally, we partition the $p$ elements of the full vector $X_{i}\in\mathbb{R}^{p}$ into blocks of size $d$: within the $b$-th block ($b=1,\dots,p/d$), we have $X_{i}^{(b)}\in\mathbb{R}^{d}$. Then, $X_{i}^{(b)}$ is generated from a multivariate distribution with a covariance matrix $\Sigma_{\text{block}}$, defined by $\Sigma_{\text{block}}=\rho\mathbf{1}\mathbf{1}^{\top}+(1-\rho)I_{d}.$ We consider two types of distributions for generating each block of features: (i) A Gaussian distribution (multivariate normal) with mean zero and covariance matrix $\Sigma_{\text{block}}$, and (ii) A multivariate $t$-distribution with degrees of freedom $\nu=5$, standardized to have covariance matrix $\Sigma_{\text{block}}$. We set $(n,p)\in\{(100,50),(200,100)\},$ and $d\in\{5,10,25,50\}$ and investigate cases where the cell size $d$ is as large as the dimension $p$.
We evaluate and compare the following four estimators for $\theta_{0}$: (i) the ordinary least squares estimator (OLS), which minimizes the squared loss function without regularization, (ii) the Ridge estimator, which minimizes the squared loss function with an $R_{0}(w)=\|\cdot\|_{2}^{2}$, (iii) the Lasso estimator which minimizes the squared loss function with an $R_{0}(w)=0.1\|\cdot\|_{1}$, and (iv) the Huber estimator, with its parameter $\eta=1.35$. For each scenario, we generate synthetic datasets repeatedly ($100$ trials) from both Gaussian and $t$-distributed designs and compute the norms of the estimation errors $\|\widehat{\theta}-\theta_{0}\|_{\Sigma}^{2}$ for each estimator.
Figure 1 and 2 show quantile-quantile plots, which plot quantiles of estimation error norms obtained under Gaussian-generated data (horizontal axis) against those obtained under $t$-distributed data (vertical axis). In all tested scenarios, points closely align with the 45-degree line, visually supporting the universality even if $d$ has the same order as $p$.
Figure 1. Q-Q plots of the estimation errors. The horizontal axis is for the Gaussian design, and the vertical axis is for the tt‑distributed design. We set $n=100,p=50$.
Figure 2. Q-Q plots of the estimation errors. The horizontal axis is for the Gaussian design, and the vertical axis is for the tt‑distributed design. We set $n=200,p=100$.